OpenStudy (anonymous):

in 7,4,1 which term is -296

OpenStudy (anonymous):

Ok d=4-7=-3 -296=7+(n-1)*-3 -303=-3(n-1) 101=(n-1) n=102

OpenStudy (saifoo.khan):

crrct.

OpenStudy (anonymous):

yes